# Ultra-low Frequency Noise External Cavity Diode Laser Systems for Quantum Applications

### v1

@@ 1 Introduction
Figure 1: Typical FN behaviour of a locked and a free-running ECDL [5] and a free-running TiSa [30]. Compared to the free running TiSa, the locked ECDL shows a smaller FN for slower Fourier frequencies. The excess noise (dark gray area) is shifted to faster FN, leading to a servo bump. The limit of FN reduction for fast Fourier frequencies is determined by the laser system and the bandwidth of the control loop.
@@
In this study, we analyze the behaviour of ULNL based on an ECDL with weak optical feedback using an additional fiber cavity at $729$ nm. This wavelength is particularly relevant for coherent manipulation of calcium ions that are utilized as optical clocks or qubits in quantum information processing and quantum simulations. These applications have strict FN requirements in the MHz range. Specifically, we investigate the influence of the light source on the system behaviour by comparing two identical laser setups, differing only in the type of the used laser diode. Our results demonstrate that the use of an FP laser diode as a source significantly improves the system’s behaviour in the form of mode stability and frequency selection. However, when utilizing a refined FP laser diode with an AR coating on the front facet [48], the system becomes much more stable in term of single-mode behaviour. These findings suggest that careful consideration of the type of laser diode used is critical for achieving stable and reliable operation of ULNL systems with weak optical feedback. This work provides valuable insights for the development and optimization of such systems for a range of applications in the field of ECDLs. In Section 2 we introduce the theory for our work, in Section 3 we describe our experimental setup including the measurement and characterisation system and in Section 4 we present the theoretical and experimental results for our ULNLs.
@@ 2 Theory
For our theoretical consideration we look at two different topics: On the one hand we describe the ULNL in terms of FN and linewidth reduction, on the other hand we focus on the mode structure of the two lasers with different types of LDs. Figure 2 presents schematically a laser based on a laser diode with two additional external cavities. The whole laser consists of four reflective surfaces resulting in three coupled cavity parts: The active medium of the laser diode (amplitude reflection coefficients $r_{1}$ and $r_{2}$ with the refractive-index-dependent cavity length $n_{\mathrm{LD}}\cdot L_{\mathrm{int}}$) forms the first cavity. The second part ($r_{\mathrm{ex1}}$ with $L_{\mathrm{ex1}}$) creates the ECDL with the frequency-selective grating [29]. The long additional fiber with an end-mirror ($r_{\mathrm{ex2}}$ with $L_{\mathrm{ex2}}$) completes the ULNL.
@@
with the linewidth enhancement factor $\alpha_{H}$ (relates phase changes to changes of the gain) [59], the separate cavity lengths $n_{\mathrm{LD}}\cdot L_{\mathrm{int}}$, $L_{\mathrm{ex1}}$ and $L_{\mathrm{ex2}}$ and the feedback ratio $\beta$, defined as the ratio between output power and feedback power. The latter is responsible for both the linewidth reduction and the stability of the system. To investigate the influence of $\beta$ on $\Delta\nu_{L,\mathrm{ULNL}}$, we can change this value in our laser experimentally over a wide range, described in Section 4. Different operation regimes for variable feedback ratios $\beta$ are described in [60, 61, 62, 33]: A too-strong feedback ratio $\beta$ leads to multiple allowed lasing mode solutions and turns single-mode behaviour to multi-mode. This effect is known as coherence collapse, which will be analysed in Section 4.
@@
Laser diode reflection coefficient rear facet $r_{\mathrm{1}}$ $0.99$
Laser diode reflection coefficient front facet $r_{\mathrm{2,FP}}$ $0.05$
$r_{\mathrm{2,AR}}$ $0.001$
@@
First external cavity reflection coefficient $r_{\mathrm{ex1}}$ $0.15$
@@
Table 1: Assumed laser parameters for the simulations.
@@
For analyzing the single longitudinal mode behaviour of an ULNL, it is necessary to focus on each cavity mode of each part of the laser system. The simplest cavity of a laser consists of two reflective surfaces in a defined distance. This concept is known as the two-mirror model. If more cavity interfaces such as partially reflective surfaces are involved, the scattering matrix formalism can be used to describe the multi-mirror laser cavity [27, 33]. This technique allows to reduce a configuration with several coupled cavities back to a two-mirror system with modified properties, without loss of generality.
To reduce the multi-mirror system to a two-mirror system, we introduce the definition of the effective amplitude reflection $r_{\mathrm{eff}}$ [63]. This complex number combines two cavity surfaces in one value:
$r_{\mathrm{eff}}=\frac{r_{i}+r_{\mathrm{j}}e^{i\omega\tau_{j}}}{1+r_{i}r_{\mathrm{j}}e^{i\omega\tau_{j}}}$ (3)
with $r_{i}$ as the amplitude reflection coefficient of the first cavity surface and $r_{j}$ as the amplitude reflection coefficient of the second cavity surface. The photon lifetime in the external cavity is $\tau_{j}=2nL_{j}/c_{0}$ and the oscillation laser frequency is $\omega=2\pi\nu$.
Figure 2: Schematic of the multi-mirror model for laser including laser diode and the two external cavities. The reduction of the system to a simple two-mirror model with $r_{1}$ and $r_{\mathrm{eff}}$ requires two steps.
For reducing the ULNL to a two-mirror model, $r_{\mathrm{eff}}$ is calculated in two steps. In the first step, the ECDL is described with $r_{i}=r_{2}$, $r_{j}=r_{\mathrm{ex1}}$ and $L_{j}=L_{\mathrm{ex1}}$, which generates an intermediate value $r_{\mathrm{eff,ECDL}}$. In the second step, the final effective reflectivity $r_{\mathrm{eff,ULNL}}$ is calculated with $r_{i}=r_{\mathrm{eff,ECDL}},r_{j}=r_{\mathrm{ex2}}$ and $L_{j}=L_{\mathrm{ex2}}$. In summary, the ULNL is reduced to the back facet amplitude reflection coefficient of the laser diode $r_{1}$ and $r_{\mathrm{eff,ULNL}}$ which includes $r_{2},r_{\mathrm{ex1}},L_{\mathrm{ex1}},r_{\mathrm{ex2}}$ and $L_{\mathrm{ex2}}$.
The distinction between an FP laser diode and an AR laser diode is the partial reflection of the laser diode front facet. For an FP laser diode the typical reflectivity of a front facet is $r_{2,FP}\approx 0.05$. To reduce this value and thus suppress the internal modes, an AR coating can be applied. This coating reduces the partial reflection down to $r_{2,AR}\approx 0.001$ [48]. Compared to diode-based systems with an FP laser diode, a reduced amplitude reflection coefficient of the front facet of the laser diode has less influence on the mode behaviour of the whole system. An AR coating strongly suppresses the inner longitudinal modes of the laser diode, while an FP laser diode adds one additional influencing surface to the cavity structure, which leads to more a complex mode structure.
@@
where $r_{i}$ and $r_{j}$ are the reflectivities of the involved cavity mirrors and the $\mathrm{FSR}=c/2nL$ are used.
Figure 3: Simulated $T(\nu)$ for an ULNL with an AR laser diode. (a) The different longitudinal mode structures for all involved cavities: The internal (blue) and the two external (orange and green) cavity modes are calculated while using the real laser parameters for the AR laser diode case. The dashed line describes the frequency selective bandpass function of the grating (GU). Note: Modeex2 is so high frequency and low in amplitude that it is barely visible. (b) Combined mode structure of an ULNL calculated with $r_{1}$ and $r_{\mathrm{eff,ULNL}}$.
@@
The complete mode structure of an ULNL is shown in Figure 3(b), which is calculated using Equation 4 with $r_{1}$, $r_{\mathrm{eff,ULNL}}$ and the total optical length of all cavity parts. Compared to Figure 3(a) it becomes obvious that the most dominant part of the system is the ECDL: The grating with a reflectivity of $r_{\mathrm{ex1}}\approx 15$ in combination with its frequency selective properties leads to single-mode laser operation (Figure 3(b) red dot). However, the additional delay line has a significant drawback: the small FSR makes the system susceptible to external influences. Small changes in the internal cavity length caused by fluctuations of electrical current or temperature, as well as variations in the external cavity length through fluctuations of air pressure or temperature shift the mode structure of each cavity part. This changes the combined transmittance profile, leading to variations in the amplitude of peaks near the main peak. If one of these peaks has higher gain than the actual emission peak, a mode jump occurs as the emission frequency jumps to the new maximum of $T$. Note that this approach to the modelling of mode behaviour of a laser resonator does not take into account any semiconductor physics effects.
@@ 3 Experimental setup
For measuring the FN spectrum we detect the power spectral density $S_{\nu}(f)$ with a delayed self-heterodyne setup. Therefore, we separate the light into two parts: One part is frequency shifted using a $40$ MHz acousto-optical modulator (AOM). The second part is delayed using a fiber with $20$ m length, which leads to a delay of $\approx 100$ ns. Both beams are overlapped on a photodiode. An oscilloscope (Teledyne LeCroy HDO6104A) detects the beat signal and a software analyses the FFT to calculate the FN spectrum of the laser for Fourier frequencies between $300$ Hz to $6$ MHz [50, 52].
@@
The connection between two independent actuating parameters (e.g. the lengths of the two external cavities) and the behaviour of the longitudinal modes of a laser can be depicted in a two dimensional plot. These so-called mode maps are generated with the measurement setup shown in Figure 4. The light is split in two paths by a BS. A wavemeter (High Finesse, WS8) with a resolution of $200$ kHz measures the frequency of the light. Simultaneously, a scanning Fabry Pérot interferometer (FPI) with an FSR $=1$ GHz is used to identify single-mode regimes of operation. Therefore the internal piezo of the FPI scans continuously the cavity length over several FSR, while the photodiode (PDFPI) behind the FPI records the transmitted intensity. When the laser is single-mode and resonant with the FPI, distinct sharp and narrow transmission peaks are detected with PDFPI. The distance between the peaks is constant. When the laser starts to run multi-mode, these main peaks first lose intensity, while additional small peaks become visible. These additional peaks originate from other excited longitudinal modes. If the additional peaks exceed a threshold of $>2$ compared to the main peak, the laser is recognized as multi-mode. This measurement is performed for any tuple of values of interest. We defined two independent variables, which are iterated through in two nested loops with specific step sizes, which are typically $\leq 1$ of the maximum allowed value. For each set of parameters a waiting period is implemented (settling time $\approx 0.5$ s) before the recording of the measured values. Each pair of values represent one point in the mode map. In single-mode regions the measured wavelength is encoded according to a color scale, while multi-mode regimes are depicted in white. The example of such a mode map is shown in Figure 6; the details will be discussed in Section 4.
@@
For laser frequency stabilization, the ULNL can be locked to the fundamental transverse mode of a plano-concave reference cavity, which has a FSR $=1.5$ GHz and a $\mathcal{F}\approx 10000$. The Pound-Drever-Hall technique [66, 67] is utilized to establish the lock. To implement this technique, an electro-optic modulator (EOM) in conjunction with a local oscillator (LO) is utilized to generate $25$ MHz sidebands. The control signal for the lock can be detected using a half-wave plate (HWP) and a PBS in reflection with a photodiode (PDref). The TOPTICA locking electronics (DLC pro with lock option, Pound-Drever-Hall module, FALC pro) are utilized to feed the control signal back to the laser system. Two actuators are employed to maintain resonance with the cavity. Fast fluctuations are corrected with the current of the laser diode, while the slow fluctuations are addressed using the piezo stack of the fiber cavity (PZTfib).
@@ 4 System comparison and discussion
In order to ensure a fair comparison between the two ULNLs, both laser systems were characterized simultaneously. Also, all components of significant importance that impact laser performance, such as the grating GU, BS, and PM fiber, were thoroughly measured and selected beforehand. This ensures that only difference between the two systems is the used laser diode type.
@@
Figure 5(b) illustrates the variation of $\Delta\nu_{L,\mathrm{ULNL}}$ as a function of $\beta$ for both ULNLs. Both systems exhibit a reduction in $\Delta\nu_{L,\mathrm{ULNL}}$ due to the presence of an additional fiber cavity. It is possible to achieve $\Delta\nu_{L,\mathrm{ULNL}}\leq 0.1$ kHz with AR and FP laser diodes. However, the system employing the FP laser diode tends to require higher values of $\beta$ to achieve the same linewidth reduction as the system with the AR laser diode with $\beta<-37$ dB. This can be attributed to the higher reflectivity of the partially reflecting front facet of the FP laser diode. It leads to increased reflection losses, resulting in reduced feedback entering the laser diode. Beyond a feedback ratio threshold of $\beta=-37$ dB, the linewidth reduction for both systems becomes equal and reaches a minimum. For $\beta<-42$ dB, the linewidth reduction decreases. Figure 5(b) also depicts the feedback ratios at which both systems undergo a transition to multi-mode behaviour: for the AR laser diode it is at $\beta>-27$ dB and for the FP laser diode it is at $\beta>-32$ dB. This effect is known as coherence collapse [60]. In the literature the optical feedback regime for a single-mode operation with narrow linewidth is known as regime III. This regime is known to be independent of feedback phase and it provides stable narrow-line laser operation. This is exactly the regime in which we operate our ULNLs. It usually spans only a small range of feedback ratios. The pioneering work of Tkach and Chraplyvy [60] provided $-45<\beta<-39$ dB as boundaries. However, these values were measured only for one type of laser diode. Further works revealed that the boundaries of regime III can depend on several parameters [61, 62, 33]: the type of laser diode, larger output powers, longer cavity lengths and smaller laser coupling factors help to increase the width or shift the boundaries of regime III. In our case we see the boundaries are shifted for both laser diodes, which could be a result of the mentioned parameters. However, the transition to the coherent collapse happens at different feedback levels for AR and FP laser diodes. In summary the ULNL with an AR laser diode achieve $\Delta\nu_{L,\mathrm{ULNL}}=0.1$ kHz in a range of $\Delta\beta\approx 13$ dB, while the ULNL with an FP laser diode achieve it within $\Delta\beta\approx 5$ dB. The broader regime III makes the AR system more stable against feedback power fluctuations.
@@
Figure 6: Measured mode maps for ULNL with an AR laser diode and an FP laser diode compared to simulated mode maps for the two ULNLs. The two independent variables are $L_{\mathrm{ex1}}$ and $L_{\mathrm{ex2}}$. The white areas represent multi-mode regimes.
@@
The measurement starts with a calibration which brings both lasers into a single-mode regime of operation. On the left side of Figure 6, the behaviour of the system employing the AR laser diode is depicted. Each mode map starts at the bottom left. First, the parameter PZTGU is fixed, while the parameter PZTfib changes in value, which leads to an increase of $L_{\mathrm{ex2}}$. The mode map demonstrates that single-mode operation is maintained throughout the whole measurement. The observed mode jumps correspond to the $\mathrm{FSR}=55$ MHz of the ULNL. Secondly, the PZTGU makes a step to the next value, the length of the first external cavity changes by around $\Delta L_{\mathrm{ex1}}=2$ nm. Correspondingly, the emission frequency of the ECDL changes by a few MHz. For this value, the same change of $\Delta L_{\mathrm{ex2}}$ leads to the same mode hop behaviour as before. In total it is possible to set every desired $\nu$ in a range of $\Delta\nu\pm 100$ MHz in the given parameter range. The same mode behaviour is observed in the simulation. Mode hops arising between longitudinal jumps of the second cavity matching its FSR are observed. The range of simulated $\Delta\nu$ over the whole parameter range is a few MHz lower as for the measured data, which we attribute to a drift of the center frequency during the measurement. Nevertheless the simulations describe the mode behaviour of the AR laser diode-based ULNL sufficiently well to predict the influence of changing cavity lengths.
@@
In contrast, the mode maps of the laser system employing the FP laser diode reveal significantly worse single-mode behaviour (Figure 6 right column). The laser regularly switches between single-mode and multi-mode operation while scanning the length of the second external cavity $L_{\mathrm{ex2}}$. The multi-mode regions are approximately twice as large as the single-mode regions. The single-mode regions are reduced to $20$ MHz tuning range and are thus significantly smaller compared to the system with the AR laser diode. This behaviour repeats for every setting of the length of the first external cavity $L_{\mathrm{ex1}}$. This represents a crucial difference compared to the AR laser diode. However, single-mode areas with limited tuning can also be found. The simulations show qualitative agreement with the measurements, but differ in the size of the multi-mode areas and the frequency deviation. This can be explained by the simple structure of the simulations, which only take into account the effect of changes in the cavity lengths, while the physics of the semiconductors dynamics is only considered empirically. The operation of a real laser is more complex than our simulation, but our simple model is well suited to qualitatively describe the mode structure of complex ECDLs.

### v2

@@ 1 Introduction
Figure 1: Typical FN behaviour of a locked and a free-running ECDL [5] and a free-running TiSa [30]. Compared to the free running TiSa, the locked ECDL shows a smaller FN for slower Fourier frequencies. The excess noise (dark grey area) is shifted to faster FN, leading to a servo bump. The limit of FN reduction for fast Fourier frequencies is determined by the laser system and the bandwidth of the control loop.
@@
In this study, we analyze the behaviour of ULNL based on an ECDL with weak optical feedback using an additional fiber cavity at $729$ nm. This wavelength is particularly relevant for coherent manipulation of calcium ions that are utilized as optical clocks or qubits in quantum information processing and quantum simulations. These applications have strict FN requirements in the MHz range. Specifically, we investigate the influence of the light source on the system behaviour by comparing two identical laser setups, differing only in the type of the used laser diode. Our results demonstrate that the use of an FP laser diode as a source significantly improves the system’s behaviour in the form of mode stability and frequency selection. However, when utilizing a refined FP laser diode with an AR coating on the front facet [48], the system becomes much more stable in terms of single-mode behaviour. These findings suggest that careful consideration of the type of laser diode used is critical for achieving stable and reliable operation of ULNL systems with weak optical feedback. This work provides valuable insights for the development and optimization of such systems for a range of applications in the field of ECDLs. In Section 2 we introduce the theory for our work, in Section 3 we describe our experimental setup including the measurement and characterisation system and in Section 4 we present the theoretical and experimental results for our ULNLs.
@@ 2 Theory
For our theoretical consideration we look at two different topics: On the one hand we describe the ULNL in terms of FN and linewidth reduction, on the other hand we focus on the mode structure of the two lasers with different types of LDs. Figure 2 presents schematically a laser based on a laser diode with two additional external cavities. The whole laser consists of four reflective surfaces resulting in three coupled cavity parts: The active medium of the laser diode (amplitude reflection coefficients $r_{1}$ and $r_{2}$ with the refractive-index-dependent cavity length $n_{\mathrm{LD}}\cdot L_{\mathrm{int}}$) forms the first cavity. The second part ($r_{\mathrm{ex1}}$ with $L_{\mathrm{ex1}}$) creates the ECDL with the frequency-selective grating [29]. The long additional fiber with an end-mirror ($r_{\mathrm{ex2}}$ with $L_{\mathrm{ex2}}$) completes the ULNL. All relevant parameters are listed in Table 1.
Figure 2: Schematic of the multi-mirror model for laser including laser diode and the two external cavities. The system can be reduced to a two-mirror laser with $r_{1}$ and $r_{\mathrm{eff}}$.
@@
with the linewidth enhancement factor $\alpha_{H}$ (relates phase changes to changes of the gain) [59], the separate cavity lengths $n_{\mathrm{LD}}\cdot L_{\mathrm{int}}$, $L_{\mathrm{ex1}}$ and $L_{\mathrm{ex2}}$ and the feedback ratio $\beta$, defined as the ratio between feedback power and output power. The latter is responsible for both the linewidth reduction and the stability of the system. To investigate the influence of $\beta$ on $\Delta\nu_{L,\mathrm{ULNL}}$, we can change this value in our laser experimentally over a wide range, described in Section 4. Different operation regimes for variable feedback ratios $\beta$ are described in [60, 61, 62, 33]: A too-strong feedback ratio $\beta$ leads to multiple allowed lasing mode solutions and turns single-mode behaviour to multi-mode. This effect is known as coherence collapse, which will be analysed in Section 4.
@@
Laser diode reflection coefficient rear facet $r_{\mathrm{1}}^{2}$ $0.99$
Laser diode reflection coefficient front facet $r_{\mathrm{2,FP}}^{2}$ $0.05$
$r_{\mathrm{2,AR}}^{2}$ $0.001$
@@
First external cavity reflection coefficient $r_{\mathrm{ex1}}^{2}$ $0.15$
@@
Table 1: Laser parameters used in the simulations.
@@
For analyzing the single longitudinal mode behaviour of an ULNL, it is necessary to focus on each cavity mode of each part of the laser system. The simplest cavity of a laser consists of two reflective surfaces in a defined distance. This concept is known as the two-mirror model. If more cavity interfaces such as partially reflective surfaces are involved, the scattering matrix formalism can be used to describe the multi-mirror laser cavity [27]. This technique allows to reduce a configuration with several coupled cavities back to a two-mirror system with modified properties, without loss of generality.
To reduce the multi-mirror system to a two-mirror system, we introduce the definition of the effective amplitude reflection $r_{\mathrm{eff}}$ [33]. This complex number combines the diode front-facet and two external cavity surfaces in one value using scattering theory assuming no losses and no dispersion:
$r_{\mathrm{eff}}=S_{11}+\frac{S_{12}S_{21}r_{\mathrm{ex}2}\,e^{-i\omega\tau_{\mathrm{ex}2}}}{1-S_{22}r_{\mathrm{ex}2}\,e^{-i\omega\tau_{\mathrm{ex}2}}}$ (3)
with the amplitude reflection coefficient of the second external cavity $r_{\mathrm{ex}2}$, the photon lifetime in the second external cavity $\tau_{\mathrm{ex}2}=2L_{\mathrm{ex}2}/c_{0}$ and the oscillation laser frequency $\omega=2\pi\nu$. The corresponding scattering coefficients $S_{ij}$ are
$S_{11}=r_{2}+\frac{t_{2}^{2}r_{\mathrm{ex1}}\,e^{-i\omega\tau_{\mathrm{ex1}}}}{1+r_{2}r_{\mathrm{ex1}}\,e^{-i\omega\tau_{\mathrm{ex1}}}},\,\,\,S_{12}=S_{21}=\frac{t_{2}t_{\mathrm{ex1}}\,e^{-\frac{1}{2}i\omega\tau_{\mathrm{ex1}}}}{1+r_{2}r_{\mathrm{ex1}}\,e^{-i\omega\tau_{\mathrm{ex1}}}},\,\,\,S_{22}=-r_{\mathrm{ex1}}-\frac{t_{\mathrm{ex1}}^{2}r_{2}\,e^{-i\omega\tau_{\mathrm{ex1}}}}{1+r_{2}r_{\mathrm{ex1}}\,e^{-i\omega\tau_{\mathrm{ex1}}}}$
with the amplitude transmission coefficients $t_{2}$ and $t_{\mathrm{ex1}}$ and $t_{i}^{2}+r_{i}^{2}=1$, the amplitude reflection coefficient of the ECDL $r_{\mathrm{ex1}}$ and the photon lifetime in the ECDL cavity $\tau_{\mathrm{ex1}}=2L_{\mathrm{ex1}}/c_{0}$. In summary, the ULNL is reduced to a laser cavity consisting of two mirrors with reflectivities $r_{1}$ and $r_{\mathrm{eff}}$.
The distinction between an FP laser diode and an AR laser diode is the partial reflection of the laser diode front facet. For an FP laser diode the typical reflectivity of a front facet is $r_{2,\mathrm{FP}}^{2}\approx 0.05$. To reduce this value and thus suppress the internal modes, an AR coating can be applied. This coating reduces the partial reflection down to $r_{2,\mathrm{AR}}^{2}\approx 0.001$ [48]. Compared to diode-based systems with an FP laser diode, a reduced amplitude reflection coefficient of the front facet of the laser diode has less influence on the mode behaviour of the whole system. An AR coating strongly suppresses the inner longitudinal modes of the laser diode, while an FP laser diode adds one additional influencing surface to the cavity structure, which leads to more a complex mode structure.
@@
where $r_{i}$ and $r_{j}$ are the reflectivities of the involved cavity mirrors and the $\mathrm{FSR}=c/2n_{i}L_{i}$ are used.
Figure 3: Simulated $T(\nu)$ for an ULNL with an AR laser diode. (a) The different longitudinal mode structures for all involved cavities: The internal (blue) and the two external (orange and green) cavity modes are calculated while using the real laser parameters for the AR laser diode case. The dashed line describes the frequency selective bandpass function of the grating (GU). Note: $\mathrm{Mode}_{\mathrm{ex2}}$ has such a high frequency that the structure cannot be resolved in the figure. (b) Combined mode structure of an ULNL calculated with $r_{1}$ and $r_{\mathrm{eff}}$.
@@
The complete mode structure of an ULNL is shown in Figure 3(b), which is calculated using Equation 4 with $r_{1}$, $r_{\mathrm{eff}}$ and the total optical length of all cavity parts. Compared to Figure 3(a) it becomes obvious that the most dominant part of the system is the ECDL: The grating with a reflectivity of $r_{\mathrm{ex1}}^{2}\approx 0.15$ in combination with its frequency selective properties leads to single-mode laser operation (Figure 3(b) red dot). However, the additional delay line has a significant drawback: the small FSR makes the system susceptible to external influences. Small changes in the internal cavity length caused by fluctuations of electrical current or temperature, as well as variations in the external cavity length through fluctuations of air pressure or temperature shift the mode structure of each cavity part. This changes the combined transmittance profile, leading to variations in the amplitude of peaks near the main peak. If one of these peaks has higher gain than the actual emission peak, a mode jump occurs as the emission frequency jumps to the new maximum of $T$. Note that this approach to the modelling of mode behaviour of a laser resonator does not take into account any semiconductor physics effects.
@@ 3 Experimental setup
For measuring the FN spectrum we detect the power spectral density $S_{\nu}(f)$ with a delayed self-heterodyne setup. Therefore, we separate the light into two parts: One part is frequency shifted using a $40$ MHz acousto-optical modulator (AOM, AA Opto-Electronic MTS40-A3-750.850). The second part is delayed using a fiber with $20$ m length, which leads to a delay of $\approx 100$ ns. Both beams are overlapped on a photodiode. An oscilloscope (Teledyne LeCroy HDO6104A) detects the beat signal and a software analyses the FFT to calculate the FN spectrum of the laser for Fourier frequencies between $300$ Hz to $6$ MHz [50, 52].
@@
The connection between two independent actuating parameters (e.g. the lengths of the two external cavities) and the behaviour of the longitudinal modes of a laser can be depicted in a two dimensional plot. These so-called mode maps are generated with the measurement setup shown in Figure 4. The light is split in two paths by a BS. A wavemeter (High Finesse, WS8) with a resolution of $200$ kHz measures the frequency of the light. Simultaneously, a scanning Fabry-Pérot interferometer (FPI) with an FSR $=1$ GHz and a $\mathcal{F}\approx 300$ is used to identify single-mode regimes of operation. Therefore the internal piezo of the FPI scans continuously the cavity length over several FSR, while the photodiode (PDFPI) behind the FPI records the transmitted intensity. When the laser is single-mode and resonant with the FPI, distinct sharp and narrow transmission peaks are detected with PDFPI. The distance between the peaks is constant. When the laser starts to run multi-mode, these main peaks first lose intensity, while additional small peaks become visible. These additional peaks originate from other excited longitudinal modes. If the additional peaks exceed a threshold of $>2$ compared to the main peak, the laser is recognized as multi-mode. This measurement is performed for any tuple of values of interest. We defined two independent variables, which are iterated through in two nested loops with specific step sizes, which are typically $\leq 1$ of the maximum allowed value. For each set of parameters a waiting period is implemented (settling time $\approx 0.5$ s) before the recording of the measured values. Each pair of values represent one point in the mode map. In single-mode regions the measured wavelength is encoded according to a color scale, while multi-mode regimes are depicted in white. The example of such a mode map is shown in Figure 6; the details will be discussed in Section 4.
@@
For laser frequency stabilization, the ULNL can be locked to the fundamental transverse mode of a plano-concave reference cavity, which has an FSR $=1.5$ GHz and a $\mathcal{F}\approx 10000$. The Pound-Drever-Hall technique [65, 66] is utilized to establish the lock. To implement this technique, an electro-optic modulator (EOM, QUBIG PM7-NIR) in conjunction with a local oscillator (LO, TOPTICA Pound-Drever-Hall module) is utilized to generate $25$ MHz sidebands. The control signal for the lock can be detected using a half-wave plate (HWP) and a PBS in reflection with a photodiode (PDref). The TOPTICA locking electronics (DLC pro with lock option, FALC pro) are utilized to feed the control signal back to the laser system. Two actuators are employed to maintain resonance with the cavity. Fast fluctuations are corrected with the current of the laser diode, while the slow fluctuations are addressed using the piezo stack of the fiber cavity (PZTfib).
@@ 4 System comparison and discussion
In order to ensure a fair comparison between the two ULNLs, both laser systems were characterized simultaneously. Also, all components of significant importance that impact laser performance, such as the grating GU, BS, and PM fiber, were thoroughly measured and selected beforehand. This ensures that the only difference between the two systems is the used laser diode type.
@@
Figure 5(b) illustrates the variation of $\Delta\nu_{L,\mathrm{ULNL}}$ as a function of $\beta$ for both ULNLs. Both systems exhibit a reduction in $\Delta\nu_{L,\mathrm{ULNL}}$ due to the presence of an additional fiber cavity. It is possible to achieve a linewidth of $\Delta\nu_{L,\mathrm{ULNL}}\leq 0.1$ kHz with AR and FP laser diodes. However, the system employing the FP laser diode tends to require higher values of $\beta$ to achieve the same linewidth reduction as the system with the AR laser diode with $\beta<-37$ dB. This can be attributed to the higher reflectivity of the partially reflecting front facet of the FP laser diode. It leads to increased reflection losses, resulting in reduced feedback entering the laser diode. Beyond a feedback ratio threshold of $\beta=-37$ dB, the linewidth reduction for both systems becomes equal and reaches a minimum. For $\beta<-42$ dB, the linewidth reduction decreases. Figure 5(b) also depicts the feedback ratios at which both systems undergo a transition to multi-mode behaviour: for the AR laser diode it is at $\beta>-27$ dB and for the FP laser diode it is at $\beta>-32$ dB. This effect is known as coherence collapse [60]. In the literature the optical feedback regime for a single-mode operation with narrow linewidth is known as regime III. This regime is known to be independent of feedback phase and it provides stable narrow-line laser operation. This is exactly the regime in which we operate our ULNLs. It usually spans only a small range of feedback ratios. The pioneering work of Tkach and Chraplyvy [60] provided $-45<\beta<-39$ dB as boundaries. However, these values were measured only for one type of laser diode. Further works revealed that the boundaries of regime III can depend on several parameters [61, 62, 33]: the type of laser diode, larger output powers, longer cavity lengths and smaller laser coupling factors help to increase the width or shift the boundaries of regime III. In our case we see the boundaries are shifted for both laser diodes, which could be a result of the mentioned parameters. However, the transition to the coherent collapse happens at different feedback levels for AR and FP laser diodes. In summary the ULNL with an AR laser diode achieve $\Delta\nu_{L,\mathrm{ULNL}}=0.1$ kHz in a range of $\Delta\beta\approx 13$ dB, while the ULNL with an FP laser diode achieve it within $\Delta\beta\approx 5$ dB. The broader regime III makes the AR system more stable against feedback power fluctuations.
@@
The measurement starts with a calibration which brings both lasers into a single-mode regime of operation. In the left column of Figure 6, the behaviour of the system employing the AR laser diode is depicted. Each mode map starts at the bottom left. First, the parameter PZTGU is fixed, while the parameter PZTfib changes in value, which leads to an increase of $L_{\mathrm{ex2}}$. The mode map demonstrates that single-mode operation is maintained throughout the whole measurement. The observed mode jumps correspond to the $\mathrm{FSR}=55$ MHz of the ULNL. Secondly, the PZTGU makes a step to the next value, the length of the first external cavity changes by around $\Delta L_{\mathrm{ex1}}=2$ nm. Correspondingly, the emission frequency of the ECDL changes by a few MHz. For this value, the same change of $\Delta L_{\mathrm{ex2}}$ leads to the same mode hop behaviour as before. In total it is possible to set every desired $\nu$ in a range of $\Delta\nu\pm 100$ MHz in the given parameter range. The same mode behaviour is observed in the simulation. Mode hops arising between longitudinal jumps of the second cavity matching its FSR are observed. The range of simulated $\Delta\nu$ over the whole parameter range is a few MHz lower than for the measured data, which we attribute to a drift of the center frequency during the measurement. Nevertheless the simulations describe the mode behaviour of the AR laser diode-based ULNL sufficiently well to predict the influence of changing cavity lengths.
Figure 6: Measured mode maps for ULNL with an AR laser diode and an FP laser diode compared to simulated mode maps for the two ULNLs based on Equation 4. The two independent variables are $L_{\mathrm{ex1}}$ and $L_{\mathrm{ex2}}$. The white areas represent multi-mode regimes.
@@
In contrast, the mode maps of the laser system employing the FP laser diode reveal significantly worse single-mode behaviour (Figure 6 right column). The laser regularly switches between single-mode and multi-mode operation while scanning the length of the second external cavity $L_{\mathrm{ex2}}$. The multi-mode regions are approximately twice as large as the single-mode regions. The single-mode regions are reduced to $20$ MHz tuning range and are thus significantly smaller compared to the system with the AR laser diode. This behaviour repeats for every setting of the length of the first external cavity $L_{\mathrm{ex1}}$. This represents a crucial difference compared to the AR laser diode. However, single-mode areas with limited tuning can also be found. The simulations show qualitative agreement with the measurements, but differ in the size of the multi-mode areas. This can be explained by the simple structure of the simulations, which only take into account the effect of changes in the cavity lengths, while the physics of the semiconductors dynamics is only considered empirically. The operation of a real laser is more complex than our simulation, but our simple model is well suited to qualitatively describe the mode structure of complex ECDLs.